### v2

@@ 4.1 Discrete uniform choice of basis over the Bloch sphere
Fig. 8 shows a $SSR$ for the $\mathrm{QPV^{\eta}_{(3,2,2)}}$ protocol obtained from the solutions of the SDP (68) using the Ncpol2sdpa package [Wit15] in Python, see [EFS] for the code. Notice that Fig. 8 shows that the security region for this protocol is greater than the $SR$ of $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$, meaning that it is more secure. However, analytical bounds on the best attack (even no loss) for these MoE games are so far only known for the BB84 game, and therefore we can not show tightness of our results beyond the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$  protocol – a gap between our best upper bounds and lower bounds remains. Numerical results from (68) show that for different arbitrary $m$, $p_{ans}$ for $p_{err}=0$ is upper bounded by $\frac{1}{m}$, which is attainable by the strategy of Alice randomly guessing $x$, measuring in this basis, broadcasting the outcome and answering if she was correct and otherwise claiming no photon.
@@
5 The $\mathrm{QPV}^{\eta,f}_{m_{\theta\varphi}}$ protocol and its security under entangled attackers
@@
For some $n\in\mathbb{N}$, consider a $2n$-bit function $f:\{0,1\}^{n}\times\{0,1\}^{n}\to[m]$. One round of the $m_{\theta\varphi}$-basis lossy-function protocol, denoted by $\mathrm{QPV}^{\eta,f}_{m_{\theta\varphi}}$, is described as in Definition 3.1 changing the range of $f$. The corresponding general attack is described as in Attack 3.2, extended by changing the range of the function $f$, and similarly for the $q$-qubit strategy.
@@
For simplicity, denote $\beta_{m_{\theta\varphi}}$ by $\beta$. We want to estimate the probability that for a randomly chosen $f$, we can find $f_{A}$ and $f_{B}$ such that the corresponding function $g$ is such that $\mathbb{P}_{x,y}[f(x,y)\in g_{m_{\theta\varphi}(f_{A}(x),f_{B}(y),\lambda)}]\geq(1-\beta)$.
$\begin{split}\mathbb{P}&[f:\exists f_{A},f_{B},\lambda\textrm{ s.t. }\mathbb{P}[x,y:f(x,y)\in g_{m_{\theta\varphi}}(f_{A}(x),f_{B}(y),\lambda)]\geq(1-\beta)]\\ &=\frac{\lvert\{f:\exists f_{A},f_{B},\lambda\textrm{s.t. }\mathbb{P}[x,y:f(x,y)\in g_{m_{\theta\varphi}}(f_{A}(x),f_{B}(y),\lambda)]\geq(1-\beta)\}\rvert}{\lvert\{f:\{0,1\}^{2n}\rightarrow[m]\}\rvert}\\ &\leq\frac{\lvert\{f:\exists f_{A},f_{B},\lambda\textrm{ s.t. }\forall x,y,f(x,y)\in g_{m_{\theta\varphi}}(f_{A}(x),f_{B}(y),\lambda)\}\rvert}{m^{2^{2n}}}\sum_{i=0}^{\beta 2^{2n}}\binom{2^{2n}}{i}(m-1)^{i}\\ &\leq\frac{1}{m^{2^{2n}}}2^{(2^{n+1}+1)k}(m-1)^{2^{2n}}(m-1)^{\beta 2^{2n}}2^{h(\beta)2^{2n}}\\ &=2^{(h(\beta)-\log m+(1+\beta)\log m-1)2^{2n}+(2^{n+1}+1)k}.\end{split}$ (76)
@@
$\mathbb{P}[\textsc{V}_{\textsc{AB}}\neq\textsc{Correct}\mid\textsc{V}_{\textsc{AB}}\neq\textsc{No photon}]\geq\beta_{m_{\theta\varphi}}[1-(\tilde{w}^{\xi}(\eta)+\Delta)].$ (78)
@@
Theorem 5.7.
If the attackers respond with probability $\eta$ and control at most $q\leq\frac{n}{2}-5$ qubits at the beginning of the $\mathrm{QPV}^{\eta,f}_{m_{\theta\varphi}}$  protocol and the prover’s error is such that $p_{err}<\beta_{m_{\theta\varphi}}[1-(\tilde{w}^{\xi}_{m_{\theta\varphi}}(\eta)+\Delta)]$, where $\Delta=0.009$, then the probability that the attackers are correct given that they respond is strictly smaller than the corresponding prover’s probability, i.e.
$\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{ Correct}\mid\textsc{V}_{\textsc{AB}}\neq\textsc{No photon}]<\mathbb{P}[\textsc{V}_{\textsc{P}}=\textsc{ Correct}\mid\textsc{V}_{\textsc{P}}\neq\textsc{No photon}]=1-p_{err},$ (79)
for $m_{\theta\varphi}$ and $\beta_{m_{\theta\varphi}}$ as in Lemma 5.5.
Theorem 5.7 is a direct consequence of Lemma 5.6. We see then that if $p_{err}<\beta_{m_{\theta\varphi}}[1-(\tilde{w}^{\xi}_{m_{\theta\varphi}}(\eta)+\Delta)]$ the $\mathrm{QPV}^{\eta,f}_{m_{\theta\varphi}}$ protocol is secure even in the presence of photon loss (more loss-tolerant than the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$ protocol) and attackers who pre-share entanglement.
The security of the sequential repetition of $\mathrm{QPV}^{\eta,f}_{m_{\theta\varphi}}$ is obtained as in the proof of Theorem 3.14 adjusting the parameters accordingly for the particular choice of $m_{\theta\varphi}$.
@@ 7 Discussion
A technical question in the QPV setting which we leave open is to obtain a tight finite statistical analysis for these protocols over multiple rounds. It seems intuitively clear that, since honest parties can achieve a better combination of error and response rate than any attacker for any round, the serial repetition of the proposed protocols should be secure – and we indeed expect this to be the case. However, attackers have a choice in what attack strategy to apply every round, and might even gain a slight amount by being adaptive, e.g., play a low-loss low-error round if the attackers had a lucky guess in a previous round. Because of this, finding the technical tools to bounds the best possible attack over many rounds is non-trivial.
The work of Johnston, Mittal, Russo, and Watrous [JMRW16] shows techniques how to handle extended non-local games, such as the monogamy-of-entanglement games, using SDPs. Because our work directly extends the earlier independent partial results of Buhrman, Schaffner, Speelman, and Zbinden [Spe16b, Chapter 5], we use a slightly-different SDP formulation. It would be very interesting to attempt to extend the results from [JMRW16] to extended non-local games where the players are allowed to return ‘loss’ with some probability – and investigate whether their SDP formulation gives equivalent results to ours.
Additionally, we note that (also for the proposed lossy protocols) an exponential gap remains between the entanglement required of the best attack known and the lower bounds we are able to prove. That is, we show security against attackers sharing a linear amount of entanglement, but we only know of an explicit attack whenever the attackers share an exponential number of qubits. More efficient attacks are known for specific functions, such as $f$ computable in logarithmic space [BFSS13] (for a routing version of the protocol, however the technique can easily be adapted), cf. [CM22], but even these take a polynomial amount of entanglement. Closing this gap remains a large and interesting open problem.
@@ Appendix A Non-local games and the NPA hierarchy
The tuple $\mathcal{S}=\{\left|\psi\right\rangle,A_{a}^{x},B_{b}^{y}\}_{x\in\mathcal{X},y\in\mathcal{Y},a\in\mathcal{A},b\in\mathcal{B}}$ is called strategy, and the set of all quantum behaviors is denoted by $\mathcal{Q}$. Abusing notation, we will denote $\mathcal{S}\in\mathcal{Q}$.
@@
    $\left|\psi\right\rangle$ is a normalized state, thus $\left\langle\psi\lvert\psi\right\rangle=1$, i.e., $G_{0,0}=1$.

### v3

@@ 4.2 Security of the $\mathrm{QPV}_{m_{\theta\varphi}}^{\eta}$ in the No-PE model
Fig. 11 shows a $SSR$ for the $\mathrm{QPV^{\eta}_{(3,2,2)}}$ protocol obtained from the solutions of the SDP (76) using the Ncpol2sdpa package [Wit15] in Python, see [EFS] for the code. Notice that Fig. 11 shows that the security region for this protocol is greater than the $SR$ of $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$, meaning that it is more secure. However, analytical bounds on the best attack (even no loss) for these MoE games are so far only known for the BB84 game, and therefore we can not show tightness of our results beyond the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$  protocol—a gap between our best upper bounds and lower bounds remains. Numerical results from (76) show that for different arbitrary $m$, $p_{ans}$ for $p_{err}=0$ is upper bounded by $\frac{1}{m}$, which is attainable by the strategy of Alice randomly guessing $x$, measuring in this basis, broadcasting the outcome and answering if she was correct and otherwise claiming no photon.
@@
5 The $\mathrm{QPV}^{\eta,f}_{m_{\theta\varphi}}$ protocol and its security against entangled attackers
@@
For some $n\in\mathbb{N}$, consider a $2n$-bit function $f:\{0,1\}^{n}\times\{0,1\}^{n}\to[m]$. One round of the $m_{\theta\varphi}$-basis lossy-function protocol, denoted by $\mathrm{QPV}^{\eta,f}_{m_{\theta\varphi}}$, is described as in Definition 3.1 changing the range of $f$. The corresponding general attack is described as in the attack to $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$, extended by changing the range of the function $f$, and similarly for the $q$-qubit strategy.
@@
For simplicity, denote $\beta_{m_{\theta\varphi}}$ by $\beta$. We want to estimate the probability that for a randomly chosen $f$, we can find $f_{A}$ and $f_{B}$ such that the corresponding function $g$ is such that $\Pr_{x,y}[f(x,y)\in g_{m_{\theta\varphi}(f_{A}(x),f_{B}(y),\lambda)}]\geq(1-\beta)$.
$\begin{split}\Pr&[f:\exists f_{A},f_{B},\lambda\textrm{ s.t. }\Pr[x,y:f(x,y)\in g_{m_{\theta\varphi}}(f_{A}(x),f_{B}(y),\lambda)]\geq(1-\beta)]\\ &=\frac{\lvert\{f:\exists f_{A},f_{B},\lambda\textrm{s.t. }\Pr[x,y:f(x,y)\in g_{m_{\theta\varphi}}(f_{A}(x),f_{B}(y),\lambda)]\geq(1-\beta)\}\rvert}{\lvert\{f:\{0,1\}^{2n}\rightarrow[m]\}\rvert}\\ &\leq\frac{\lvert\{f:\exists f_{A},f_{B},\lambda\textrm{ s.t. }\forall x,y,f(x,y)\in g_{m_{\theta\varphi}}(f_{A}(x),f_{B}(y),\lambda)\}\rvert}{m^{2^{2n}}}\sum_{i=0}^{\beta 2^{2n}}\binom{2^{2n}}{i}(m-1)^{i}\\ &\leq\frac{1}{m^{2^{2n}}}2^{(2^{n+1}+1)k}(m-1)^{2^{2n}}(m-1)^{\beta 2^{2n}}2^{h(\beta)2^{2n}}\\ &=2^{(h(\beta)-\log m+(1+\beta)\log m-1)2^{2n}+(2^{n+1}+1)k}.\end{split}$ (84)
@@
$\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\textsc{c}\right]\leq\eta(1-\beta_{m_{\theta\varphi}}[1-(\tilde{w}^{\epsilon}(\eta)+\Delta)]).$ (86)
@@
We see then that if $p_{err}<\beta_{m_{\theta\varphi}}[1-(\tilde{w}^{\epsilon}_{m_{\theta\varphi}}(\eta)+\Delta)]$, attackers who pre-share entanglement cannot answer ‘Correct’ with sufficiently high probability even in the presence of photon loss (more loss-tolerant than the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$ protocol). The security of the sequential repetition of $\mathrm{QPV}^{\eta,f}_{m_{\theta\varphi}}$ is obtained as in the proof of Theorem 3.13 adjusting the parameters accordingly for the particular choice of $m_{\theta\varphi}$.
@@ 7 Discussion
A technical question in the QPV setting which we leave open is to obtain a tight finite statistical analysis for these protocols over multiple rounds. It seems intuitively clear that, since honest parties can achieve a better combination of error and response rate than any attacker for any round, the serial repetition of the proposed protocols should be secure—and we indeed expect this to be the case. However, attackers have a choice in what attack strategy to apply every round, and might even gain a slight amount by being adaptive, e.g. play a low-loss low-error round if the attackers had a lucky guess in a previous round. Because of this, finding the technical tools to bounds the best possible attack over many rounds is non-trivial.
The work of Johnston, Mittal, Russo, and Watrous [JMRW16] shows techniques how to handle extended non-local games, such as the monogamy-of-entanglement games, using SDPs. Because our work directly extends the earlier independent partial results of Buhrman, Schaffner, Speelman, and Zbinden [Spe16b, Chapter 5], we use a slightly-different SDP formulation. It would be very interesting to attempt to extend the results from [JMRW16] to extended non-local games where the players are allowed to return ‘loss’ with some probability—and investigate whether their SDP formulation gives equivalent results to ours.
Additionally, we note that (also for the proposed lossy protocols) an exponential gap remains between the entanglement required of the best attack known and the lower bounds we are able to prove. That is, we show security against attackers sharing a linear amount of entanglement, but we only know of an explicit attack whenever the attackers share an exponential number of qubits. More efficient attacks are known for specific functions, such as $f$ computable in logarithmic space [BFSS13] (for a routing version of the protocol, however the technique can easily be adapted), cf. [CM23], but even these take a polynomial amount of entanglement. Closing this gap remains a large and interesting open problem.
@@ Appendix A Non-local games and the NPA hierarchy
The tuple $\mathbf{S}=\{\left|\psi\right\rangle,A_{a}^{x},B_{b}^{y}\}_{x\in\mathcal{X},y\in\mathcal{Y},a\in\mathcal{A},b\in\mathcal{B}}$ is called strategy, and the set of all quantum behaviors is denoted by $\mathcal{Q}$. Abusing notation, we will denote $\mathbf{S}\in\mathcal{Q}$.
@@
    $\left|\psi\right\rangle$ is a normalized state, thus $\left\langle\psi\lvert\psi\right\rangle=1$, i.e. $G_{0,0}=1$.